# Generalized Measures of Anticipation and Responsivity in Online Language Processing

### v1

@@ 1 Introduction
Estimates of surprisal and entropy based on neural language models have demonstrated significant predictive capacity for a wide variety of neural and behavioral data collected using self-paced and eye-tracked reading experiments (Goodkind and Bicknell, 2018; Wilcox et al., 2023), as well as EEG (Merkx and Frank, 2021; Michaelov et al., 2024), fMRI (Shain et al., 2020; Bhattasali and Resnik, 2021), and ECoG imaging (Schrimpf et al., 2021), along with explicit grammaticality and acceptability ratings Lau et al. (2017); Wallbridge et al. (2022). Despite surprisal and entropy’s empirical success, there is increasing interest in defining and evaluating alternative measures. Examples include measures designed to disentangle different dimensions—e.g., lexical versus syntactic—of uncertainty (Roark et al., 2009; Arehalli et al., 2022; Giulianelli et al., 2023) or to quantify uncertainty over spans larger than a single unit (Aurnhammer and Frank, 2019; Giulianelli et al., 2024). While estimating surprisal and entropy from neural language models is convenient, experimenters should chart new territory by designing their own measures, rethinking and enhancing established information-theoretic quantities to capture overlooked aspects of online language processing.
@@ 2.2 Generalizing Surprisal
We call ${\mathrm{f}}$ the warping function because it applies a transformation to the expected score and thus controls the shape of the distribution of generalized surprisal values for a given score distribution. It is useful to think of the warping function as characterizing the relationship between prediction accuracy and a certain construct or measurement of interest. For instance, in § 6.2.1, we show how the same notion of prediction accuracy captured by surprisal’s scoring function, ${\mathbbm{1}\{{\boldsymbol{w}}\preceq{\boldsymbol{v}}\}}$, is in a nearly linear relationship with human cloze probabilities yet in a logarithmic relationship with human predictability ratings. Much psycholinguistic research aiming to establish the functional relationship between surprisal and processing difficulty (Smith and Levy, 2013; Brothers and Kuperberg, 2021; Wilcox et al., 2023; Shain et al., 2024, inter alia) can be seen as testing different hypotheses about the workings of online language processing by instantiating them through varying warping functions.
@@ 3.1 Responsive Measures
Whilst the use of a binary scoring function follows naturally from the surprisal model of cognitive cost, it results in a relatively simplistic notion of prediction accuracy which conflates different aspects of predictive accuracy and does not take into account the communicative equivalence of predictions and observations. In the information value model of cost Giulianelli et al. (2023, 2024), instead, prediction accuracy is a continuous score that quantifies the representational distance between predictions and observations. For instance, if the predicted continuation is syntactically different but semantically equivalent to the observed next unit, this results in high syntactic and low semantic information value.
@@ 4 Experimental Setup
The information value models introduced in § 3.1 rely on a function ${\mathrm{d}}_{{\boldsymbol{c}}}$ to quantify the distance between two strings ${\boldsymbol{v}}$ and ${\boldsymbol{w}}$ in the context of ${\boldsymbol{c}}$. Here, we use mean-pooled non-contextual (i.e., layer 0) representation from GPT-2 Small and GPT-Neo 125M as column vectors ${\mathrm{r}}({\boldsymbol{v}}),{\mathrm{r}}({\boldsymbol{w}})\in\mathbb{R}^{{D}}$, where ${D}$ is the dimensionality of the embeddings, and calculate the cosine distance ${\mathrm{d}}_{{\boldsymbol{c}}}({\boldsymbol{v}},{\boldsymbol{w}})=1-\frac{{\mathrm{r}}({\boldsymbol{v}})^{\intercal}{\mathrm{r}}({\boldsymbol{w}})}{\|{\mathrm{r}}({\boldsymbol{v}})\|_{2}\|{\mathrm{r}}({\boldsymbol{w}})\|_{2}}$. There exist several other choices for distance and representation functions; see Giulianelli et al. (2023, 2024) and Meister et al. (2024) for some examples.
@@ 6.1 Evaluation
For the experiments on responsive measures (§ 6.2.1), all regressors include three baseline predictors: the length of the target string ${\boldsymbol{w}}^{({m})}$, its frequency, and the length of the context string ${\boldsymbol{c}}^{({m})}$.777 The length of the target string is measured in characters, the length of the context in words. Frequencies are extracted from the SUBTLEXus (Brysbaert et al., 2012). We call this the default baseline. For the experiments on anticipatory measures (§ 6.2.2), we use an additional baseline. Because expected next-symbol surprisal has proven to be most effective as a predictor when used in conjunction with surprisal (Pimentel et al., 2023), we compare the other, as yet untested, anticipatory measures against a baseline that includes both surprisal and expected next-symbol surprisal, along with the three baseline predictors mentioned above (the combined baseline). In the target regressor, expected next-symbol surprisal is then replaced with another anticipatory target predictor, allowing us to assess the boost, or decrease, in predictive power that results from this substitution.888 § D.2 also reports results for the default baseline.
In our analysis, we only use linear regression so that the functionalauthor=Ryan,color=green!30,size=,fancyline,caption=,]I would have a nice footnote relating the warpoing function to, say, GAMs here. Mario: How about this? relationship between the scoring function and the psycholinguistic variable can be unambiguously expressed via the warping function.999 Others (e.g., Smith and Levy, 2008, 2013; Goodkind and Bicknell, 2018; Brothers and Kuperberg, 2021; Wilcox et al., 2023) have tried to learn the form of this relationship from data, using generalized additive models Wood (2004, 2017). The regressors are fit with the ordinary least squares method and evaluated on a held-out test set. We quantify the predictive power of a measure as the difference in the coefficient of determination ${R^{2}}$ of the target regressor and the baseline regressor, denoted as ${\Delta_{{R^{2}}}}$. The statistical significance of a measure’s ${\Delta_{{R^{2}}}}$ is assessed via 10-fold cross-validation and permutation tests. A full description of our analysis procedure is given in § D.1.
@@ 7 Conclusion
We introduced a generalization over classic information-theoretic measures of predictive uncertainty in online language processing. Our generalized surprisal framework subsumes both responsive and anticipatory measures, including established special cases but providing a vocabulary and the formal tools for experimenters to design new measures and explain psycholinguistic data of interest.
@@ A.2 Aggregating Multi-Token Estimates
When a word is composed of multiple subword tokens, token-level estimates are aggregated. For probability-based measures, we naturally multiply token-level estimates following the chain rule. For surprisal and information value, we sum token-level estimates (which, for surprisal, is equivalent to multiplying token probabilities).

### v2

@@ 1 Introduction
Estimates of surprisal and entropy based on neural language models have demonstrated significant predictive capacity for a wide variety of neural and behavioral data collected using self-paced and eye-tracked reading experiments (Goodkind and Bicknell, 2018; Wilcox et al., 2023), as well as EEG (Merkx and Frank, 2021; Michaelov et al., 2024), fMRI (Shain et al., 2020; Bhattasali and Resnik, 2021), and ECoG imaging (Schrimpf et al., 2021), along with explicit grammaticality and acceptability ratings Lau et al. (2017); Wallbridge et al. (2022). Despite surprisal and entropy’s empirical success, there is increasing interest in defining and evaluating alternative measures. Examples include measures designed to disentangle different dimensions—e.g., lexical versus syntactic—of uncertainty (Roark et al., 2009; Arehalli et al., 2022; Giulianelli et al., 2023) or to quantify uncertainty over spans larger than a single unit (Aurnhammer and Frank, 2019; Giulianelli et al., 2024b). While estimating surprisal and entropy from neural language models is convenient, experimenters should chart new territory by designing their own measures, rethinking and enhancing established information-theoretic quantities to capture overlooked aspects of online language processing.
@@ 2.2 Generalizing Surprisal
We call ${\mathrm{f}}$ the warping function because it applies a transformation to the expected score and thus controls the shape of the distribution of generalized surprisal values for a given score distribution. It is useful to think of the warping function as characterizing the relationship between prediction accuracy and a certain construct or measurement of interest. For instance, in § 6.2.1, we show how the same notionauthor=Ryan,color=green!30,size=,fancyline,caption=,]Why is it binary? I think it’s still in $[0,1]$ rather than $\{0,1\}$.  Mario: It is in $[0,1]$ in expectation, but the scoring function itself is binary, do you agree? of prediction accuracy captured by surprisal’s scoring function, ${\mathbbm{1}\{{\boldsymbol{w}}\preceq{\boldsymbol{v}}\}}$, is in a nearly linear relationship with human cloze probabilities yet in a logarithmic relationship with human predictability ratings. Much psycholinguistic research aiming to establish the functional relationship between surprisal and processing difficulty (Smith and Levy, 2013; Brothers and Kuperberg, 2021; Wilcox et al., 2023; Shain et al., 2024, inter alia) can be seen as testing different hypotheses about the workings of online language processing by instantiating them through varying warping functions.
@@ 3.1 Responsive Measures
Whilst the use of a binary scoring function follows naturally from the surprisal model of cognitive cost, it results in a relatively simplistic notion of prediction accuracy which conflates different aspects of predictive accuracy and does not take into account the communicative equivalence of predictions and observations. In the information value model of cost Giulianelli et al. (2023, 2024b), instead, prediction accuracy is a continuous score that quantifies the representational distance between predictions and observations. For instance, if the predicted continuation is syntactically different but semantically equivalent to the observed next unit, this results in high syntactic and low semantic information value.
@@ 4 Experimental Setup
The information value models introduced in § 3.1 rely on a function ${\mathrm{d}}_{{\boldsymbol{c}}}$ to quantify the distance between two strings ${\boldsymbol{v}}$ and ${\boldsymbol{w}}$ in the context of ${\boldsymbol{c}}$. Here, we use mean-pooled non-contextual (i.e., layer 0) representation from GPT-2 Small and GPT-Neo 125M as column vectors ${\mathrm{r}}({\boldsymbol{v}}),{\mathrm{r}}({\boldsymbol{w}})\in\mathbb{R}^{{D}}$, where ${D}$ is the dimensionality of the representations, and calculate the cosine distance ${\mathrm{d}}_{{\boldsymbol{c}}}({\boldsymbol{v}},{\boldsymbol{w}})=1-\frac{{\mathrm{r}}({\boldsymbol{v}})^{\intercal}{\mathrm{r}}({\boldsymbol{w}})}{\|{\mathrm{r}}({\boldsymbol{v}})\|_{2}\|{\mathrm{r}}({\boldsymbol{w}})\|_{2}}$. There exist several other choices for distance and representation functions; see Giulianelli et al. (2023, 2024b) and Meister et al. (2024) for some examples.
@@ 6.1 Evaluation
For the experiments on responsive measures (§ 6.2.1), all regressors include three baseline predictors: the length of the target string ${\boldsymbol{w}}^{({m})}$, its frequency, and the length of the context string ${\boldsymbol{c}}^{({m})}$. [^7] The length of the target string is measured in characters, the length of the context in words. Frequencies are extracted from the SUBTLEXus (Brysbaert et al., 2012). We call this the default baseline. For the experiments on anticipatory measures (§ 6.2.2), we use an additional baseline. Because expected next-symbol surprisal has proven to be most effective as a predictor when used in conjunction with surprisal (Pimentel et al., 2023), we compare the other, as yet untested, anticipatory measures against a baseline that includes both surprisal and expected next-symbol surprisal, along with the three baseline predictors mentioned above (the combined baseline). In the target regressor, expected next-symbol surprisal is then replaced with another anticipatory target predictor, allowing us to assess the boost, or decrease, in predictive power that results from this substitution.888 § D.2 also reports results for the default baseline.
In our analysis, we only use linear regression so that the functionalauthor=Ryan,color=green!30,size=,fancyline,caption=,]I would have a nice footnote relating the warpoing function to, say, GAMs here. Mario: How about this? relationship between the scoring function and the psycholinguistic variable can be unambiguously expressed via the warping function.999 Others (e.g., Smith and Levy, 2008, 2013; Goodkind and Bicknell, 2018; Brothers and Kuperberg, 2021; Wilcox et al., 2023) have tried to learn the form of this relationship from data, using generalized additive models Wood (2004, 2017). The regressors are fit with the ordinary least squares method and evaluated on a held-out test set. We quantify the predictive power of a measure as the difference in the coefficient of determination ${R^{2}}$ of the target regressor and the baseline regressor, denoted as ${\Delta_{{R^{2}}}}$.
The statistical significance of a measure’s ${\Delta_{{R^{2}}}}$ is assessed via 10-fold cross-validation and permutation tests. A full description of our analysis procedure is given in § D.1.
@@ 7 Conclusion
We introduced a generalization over classic information-theoretic measures of predictive uncertainty in online language processing. Our generalized surprisal framework subsumes both responsive and anticipatory measures, including established special cases, but providing a vocabulary and the formal tools for experimenters to design new measures and explain psycholinguistic data of interest.
@@ A.2 Aggregating Multi-Token Estimates
When a word is composed of multiple subword tokens, token-level estimates are aggregated. For probability-based measures, we naturally multiply token-level estimates following the chain rule. For surprisal and information value, we sum token-level estimates (which, for surprisal, is equivalent to multiplying token probabilities). See Giulianelli et al. (2024a) for a discussion on the proper treatment of tokenization in computational psycholinguistics.